OpenStudy (anonymous):

a) Calculate the following (geometric) sums and series, respectively. i) \(\Large \sum_{n=-2}^{10}2^{n}\)

OpenStudy (campbell_st):

so the series is \[\frac{1}{4} + \frac{1}{2} + 1 + 2 + 4 ....\] a = 1/4 r = 2 and n = 10 then \[s _{n} = \frac{a(r^n - 1)}{r -1} \] just substitute and the evaluate

OpenStudy (campbell_st):

oops n = 12 as you start at -2 and finish at 10

OpenStudy (anonymous):

ok.. thank you @campbell_st